OpenStudy (anonymous):

36c^2-84c+49... factor the special product.

OpenStudy (saifoo.khan):

\[(7-6 c)(7-6 c)\]

OpenStudy (anonymous):

maybe this will stump you lol.... \[t^2+(2/5)t+(1/24)\]

OpenStudy (anonymous):

last number is 25 sorry

OpenStudy (saifoo.khan):

this thing cannot be factored.

OpenStudy (anonymous):

hmmm its is one of my problems maybe i put it wrong cause it has an answer cause i tried the answer unfactorable

OpenStudy (anonymous):

\[t^2+\left(\begin{matrix}2 \\ 5\end{matrix}\right)t+\left(\begin{matrix}1 \\ 25\end{matrix}\right)\]